# Neutrinoless Double Beta Decay in Light of JUNO First Data

### v1

@@ 1 Introduction
The baryon asymmetry of the Universe, namely that almost no anti-matter exists in Nature [Cohen98, Canetti12], is one of the main problems in contemporary astroparticle physics and cosmology. An inflationary phase at the beginning of our Universe dilutes all pre-existing matter [Guth81] which requires a mechanism to understand the observed baryon asymmetry. The Sakharov conditions [Sakharov:1967dj] summarize the essential ingredients which must be fulfilled. One very well motivated mechanism is leptogenesis [Fukugita86] which is typically related to Majorana neutrino masses with additional CP phases emerging by the seesaw mechanism [Fritzsch:1975sr, Minkowski:1977sc, Yanagida:1979as, Gell-Mann:1979vob, Glashow:1979nm, Mohapatra:1979ia, Konetschny:1977bn, Magg:1980ut, Schechter:1981cv, Cheng:1980qt, Lazarides:1980nt, Mohapatra:1980yp, Foot:1988aq, Ma:1998dn].
Verifying the Majorana nature of neutrinos is therefore very important for justifying the vanilla leptogenesis mechanism. Although the ultra-violet physics at high energy is difficult to observe with current technology, with two exceptions of cosmic string [Dror:2019syi] and cosmological collider [Cui:2021iie], the nature inherited by light active neutrinos can manifest itself in the neutrinoless double beta decay ($0\nu 2\beta$) process where a mother nuclei decays into a daughter nuclei plus two electrons, $^{Z}_{A}N\rightarrow{}^{Z+2}_{A}N+2e^{-}$ [Dolinski:2019nrj, Shergold:2021evs, Agostini:2022zub, Cirigliano:2022oqy]. With two electrons appearing in the final state, such process indicates that the lepton number is violated by two. In the minimal scenario, the $0\nu 2\beta$ decay can be induced by an effective Majorana mass $m_{ee}$ of the light active neutrinos in the electron flavor. Reciprocally, a lepton number violating process would finally contribute to a light Majorana mass [Schechter82].
@@
where $G^{0\nu}$ is the phase space factor [Kotila:2012zza, Kotila:2013gea, Mirea:2014dza, Neacsu:2015uja] associated with the decay kinematics and atomic physics, $M^{0\nu}$ is the nuclear matrix element (NME), and $\langle m_{ee}\rangle$ is the effective electron neutrino mass from particle physics. Of these three parts, the phase space factor can be calculated very precisely while the NME suffers a lot from theoretical uncertainties [Agostini:2022zub]. The NME calculations involve complex nuclear many-body physics and depend on the nuclear structure models. There are multiple theoretical approaches such as the quasiparticle random phase approximation (QRPA) [qrpa1, qrpa2, qrpa3, qrpa4, qrpa5], interacting shell mode (ISM) [shell1, shell2, shell3, shell4], energy density functional methods (EDF) [edf1, edf2, edf3], interacting boson model (IBM) [ibm1, ibm2], and ab initio techniques [abinitio1, abinitio2, abinitio3, abinitio4, abinitio5, abinitio6, abinitio7]. Their predictions typically differ by a factor of $2\sim 3$, representing the main source of systematic uncertainty from the nuclear physics side [Agostini:2022zub].
On the other hand, the particle physics side has improved quite significantly. The improvement comes from neutrino oscillation experiments, beta decay spectrum measurement around the end point, and cosmology. Especially, there is complementarity between the reactor neutrino oscillation experiments and $0\nu 2\beta$ decay [Pascoli:2005zb, Choubey:2005rq, Lindner:2005kr, Dueck:2011hu, GE1, GE2, Ge:2019ldu] since both involves only the electron flavor and share the same set of neutrino oscillation parameters. To be exact, only two mixing angles, the reactor angle $\theta_{r}\equiv\theta_{13}$ and the solar angle $\theta_{a}\equiv\theta_{12}$ can appear in the effective Majorana mass $m_{ee}$ for $0\nu 2\beta$ decay and the disappearence probability $P_{ee}$ of reactor neutrino oscillation. In addition, the solar mass squared difference $\Delta m^{2}_{s}\equiv\Delta m^{2}_{21}$ and the atmospheric one $\Delta m^{2}_{a}\equiv\Delta m^{2}_{31}$ can also enter. It is interesting to see that these four oscillation parameters can be precisely measured with the combination of Daya Bay and JUNO [Ge:2012wj, JUNO:2015zny]. Especially, the recent global fits [NuFIT6] and JUNO first data [JUNO1st] have significantly reduced the uncertainty of the solar angle $\theta_{s}$. In addition, the KATRIN experiment provides the best measurement of the beta decay spectrum to put a $\sum_{i}m_{i}\leq 0.45$ eV limit at 90% C.L. (confidence level) [KATRIN2022, KATRIN2024]. The most stringent bound on the absolute neutrino mass comes from the recent DESI BAO data [DESI2024, DESI2024b], plus other existing observations such as CMB [Planck:2018vyg].
@@ 2 Effective Majorana Mass for $0\nu 2\beta$ Decay
where $m_{i}$ are the neutrino mass eigenvalues and $U_{ei}$ are the PMNS mixing matrix elements. We follow the PMNS matrix parametrization [PDG24],
@@
A particularly illuminating perspective emerges when treating $\langle m_{ee}\rangle$ as a vector sum in the complex plane which can give a geometrical view [Vissani:1999tu, Xing:2014yka, GE2]. We can decompose it into three contributions,
@@
The behavior of $\langle m_{ee}\rangle$ significantly depends on the ordering of the neutrino mass, with fundamentally different phenomenological implications for each scenario. The main distinction arises from the lightest mass, $m_{1}$ for NO and $m_{3}$ for IO. This difference has direct consequences for the allowed range of $\langle m_{ee}\rangle$, particularly in the limit where the lightest mass is smaller than $0.1$ eV. Note that distinguishing the mass ordering is the major physics goal of the JUNO experiment [JUNO:2015zny].
@@
The interplay between the neutrino oscillation parameters and the predictions for neutrinoless double-beta decay has been extensively studied in the literature, with detailed analysis of how experimental uncertainties propagate into the allowed ranges for the effective Majorana mass $\langle m_{ee}\rangle$ [Pascoli:2002xq, Pascoli:2005zb, Choubey:2005rq, Lindner:2005kr, Dueck:2011hu, GE1]. In this section, we perform an updated analysis of [GE1] by incorporating the latest global fit results from neutrino oscillation experiments [NuFIT6] and more crucially the first measurement of solar oscillation parameters $\sin^{2}\theta_{s}$ and $\Delta m^{2}_{s}$ from the JUNO experiment [JUNO1st]. Our goal is to quantify how the improved determination of the solar mixing angle $\theta_{s}$ from the JUNO initial 59.1-day dataset impacts the theoretical predictions of $\langle m_{ee}\rangle$.
@@
Although both mass squared differences are involved in the electron flavor effective mass $\langle m_{ee}\rangle$, the limits are dominated by the atmohsperic one $\Delta m^{2}_{a}$ with much larger value. Further, the upper limit $\langle m_{ee}\rangle^{\rm IO}_{\rm max}$ even becomes independent of the solar angle $\theta_{s}$. Since both the atmospheric mass squared difference $\Delta m^{2}_{a}$ and the reactor angle $\theta_{r}$ have been measured quite precisely with just roughly 1% error [NuFIT6], there is almost no uncertainty in the upper limit for IO. This has been clearly shown in Fig. 1 as the filled regions for IO have roughly the same upper boundary.
@@
For comparison, the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ depends on not just $\Delta m^{2}_{a}$ and the reactor angle $\theta_{r}$ but also the solar angle $\theta_{s}$. Although $\theta_{s}$ has also been measured very precisely with only 2% uncertainty [NuFIT6], the combination $\cos 2\theta_{s}$ that appears in $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ can vary by 1/3 within the $3\,\sigma$ range. Even in the left panel of Fig. 1 with logarithmic scale, the effect of improving the solar angle measurement with JUNO can be clearly seen. Since the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ corresponds to the maximal value of half-life time $T^{0\nu}_{1/2}$ according to Eq. (1.1), such uncertainty would have significant consequences when interpreting the experimental data. Any improvement in the determination of $\theta_{s}$ would significantly reduce the uncertainty on the lower bound of $\langle m_{ee}\rangle$ for IO [Pascoli:2002xq, GE1].
@@
These two limits contain two mass squared differences and two mixing angles. Since the reactor angle $\theta_{r}$ is not large, its consine is roughly $c_{r}\approx 1$. The relative size between the two terms in Eq. (3.3) is then determined by comparing the ratios $\Delta m^{2}_{s}/\Delta m^{2}_{a}$ and $s^{2}_{r}/s^{2}_{s}$. According to the global fit [NuFIT6], $\sqrt{\Delta m^{2}_{s}/\Delta m^{2}_{a}}\approx 0.17$ and $s^{2}_{r}/s^{2}_{s}\approx 0.07$ which means the first term $\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}$ dominates. Note that the solar angle $s^{2}_{s}$ appears in the first term. The current $3\,\sigma$ variation of $s^{2}_{s}$ can be as large as 23% [NuFIT6]. Although the solar mass squared difference $\Delta m^{2}_{s}$ itself varies by $\sim 15\%$ at $3\,\sigma$, its contribution to $\langle m_{ee}\rangle$ is only $\sim 7.5\%$. So $\theta_{s}$ is the dominant source of uncertainty for NO. With all factors combined, the total effect of those uncertainties from the current data can be clearly seen in the orange bands of the left and right panels of Fig. 1. The difference is much more visible for the lower limits. This is because the variation in the first term $\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}$ actually affects the central values as well and the whole NO branch in the limit of vanishing lightest mass moves vertically. While the lower limit $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$ is different between the two terms in Eq. (3.3), the upper one $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$ is a sum with larger value. Then the relative variation is much smaller for the upper limit.
@@
To make the influence of $\theta_{s}$ on the effective mass $\langle m_{ee}\rangle$ more visible, we have included two linear scale plots for IO and NO as the middle and right panels of Fig. 1. From this figure, it is also possible to infer from the orange bands both the minimum and maximum values of the effective mass for each mass ordering in the vanishing lightest mass scenario. Explicitly, their $3\,\sigma$ intervals are $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(15.72\sim 22.28)\,\text{meV}$, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}=(1.188\sim 1.828)\,\text{meV}$, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}=(3.404\sim 4.028)\,\text{meV}$, resulting from the propagation of the oscillation parameter uncertainties. While these uncertainties appear modest in $\langle m_{ee}\rangle$, their influence on the $0\nu 2\beta$ experiments is much more significant [GE1]. Especially, the half-life time scales inversely with the effective mass squared, $T^{0\nu}_{1/2}\propto 1/|\langle m_{ee}\rangle|^{2}$ according to Eq. (1.1). The variation in $\langle m_{ee}\rangle$ would get amplified when propagating to the half-life time. Further, to reach the required sensitivity on $T^{0\nu}_{1/2}$ the expected experimental exposure scales [Avignone:2005cs], $M\times t\propto T^{0\nu}_{1/2}$ where $M$ is the detector target mass and $t$ is the running time, with the half-life time squared. Finally, the experimental exposure can differ by a factor of 4.3 for IO and 2.5 for NO. The improvement on the uncertainty from the particle physics side by JUNO is expected to significantly reduce the uncertainty in the experimental design and data interpretation.
@@
Following the first data release from JUNO, which reported 59.1 days of data taking [JUNO1st], the precision on the solar angle $\theta_{s}$ and the solar mass splitting $\Delta m^{2}_{s}$ has been significantly improved to $\sin^{2}\theta_{s}=0.3092\pm 0.0087$ and $\Delta m^{2}_{s}=(7.50\pm 0.12)\times 10^{-5}\,$eV2. This corresponds to a reduction of uncertainty by approximately a factor of $0.741$ at $3\,\sigma$ than the global fit results from NuFIT 6.0 [NuFIT6]. According to Eq. (3.2), this translates to around 22% reduction in the uncertainty of $\langle m_{ee}\rangle^{\rm IO}_{\rm min}\propto c^{2}_{s}-s^{2}_{s}$ at $3\,\sigma$. This improvement was achieved through JUNO’s unique capability to observe the low-frequency mode of the reactor antineutrino oscillation [Ge:2012wj],
@@
For comparison, Fig. 1 shows four uncertainty scenarios. First, the theoretical prediction with perfect knowledge of oscillation parameters (red) that represents the ultimate goal of mitigating the uncertainty from the particle physics side to a negligible extend [GE1]. Second, the current $3\,\sigma$ range with the pre-JUNO global fits (orange) that shows what other experiments have achieved altogether [NuFIT6]. Between these two cases, there is a huge space which is exactly the existing uncertainty from particle physics. In the middle, we have show two uncertainty scenario from the JUNO experiment. The blue band shows the reduced uncertainty with the JUNO first data release [JUNO1st]. The comparison between the orange (pre-JUNO) and blue (post-JUNO) bands clearly illustrates the impact of the improved solar angle measurement by JUNO with a noticeable reduction in the width of the allowed regions for both mass orderings. The green one corresponds to the expected sensitivity with 6 years of data collection [JUNO:2015zny].
@@
For comparison, we have also shown those expected sensitivities from $0\nu 2\beta$ decay experiments as horizontal lines. The most stringent current bounds come from KamLAND-Zen (136Xe), yielding $T_{1/2}^{0\nu}>2.3\times 10^{26}$ yr at $90\%$ C.L., which translates to $\langle m_{ee}\rangle<(36\sim 156)$ meV depending on the nuclear matrix elements (NME) [KamLAND-Zen:2024eml, KamLAND-Zen:2022tow]. This bound is indicated as the hashed region with solid boundaries in the left and middle panels of Fig. 1. The current experiments are beginning to probe the IO parameter space, particularly in the region where $m_{\text{lightest}}\gtrsim 10$ meV. The next-generation $0\nu\beta\beta$ experiments, including nEXO [nEXO1, nEXO2], LEGEND-1000 [LEGEND, LEGEND:2021bnm], and CUPID [CUPID1, CUPID2], are projected to reach sensitivities of $\langle m_{ee}\rangle<(5\sim 12)$ meV, which would fully cover the IO prediction band and begin to probe the NO region.
@@
Fig. 2: One-dimensional $\Delta\chi^{2}$ for the $0\nu 2\beta$ effective mas limits $\langle m_{ee}\rangle^{\text{IO}}_{\text{min}}$ (left), $\langle m_{ee}\rangle^{\text{NO}}_{\text{min}}$ (center), and $\langle m_{ee}\rangle^{\text{NO}}_{\text{max}}$ (right) according to the neutrino oscillation global fit (blue) [NuFIT6] or the combined results with both global fit and JUNO (red) [JUNO1st].
@@
To make the influence of the JUNO first data [JUNO1st] on the $0\nu 2\beta$ effective mass limits more explicit, Fig. 2 shows the one-dimensional $\Delta\chi^{2}$ profiles for $\langle m_{ee}\rangle^{\text{IO}}_{\text{min}}$, $\langle m_{ee}\rangle^{\text{NO}}_{\text{min}}$, and $\langle m_{ee}\rangle^{\text{NO}}_{\text{max}}$ with vanishing lightest mass ($m_{\text{lightest}}\to 0$). In this limit, the residual uncertainty is contributed by the oscillation parameters $\Delta m^{2}_{a}$, $\Delta m^{2}_{s}$, $\theta_{s}$, and $\theta_{r}$, as discussed before. The gray dashed horizontal lines indicate the $1\,\sigma$, $2\,\sigma$, and $3\,\sigma$ confidence levels ($\Delta\chi^{2}=1$, $4$, and $9$, respectively). For comparison, the blue curves represent the NuFIT 6.0 global fit of neutrino oscillation experiments [NuFIT6] while the red curves show the improvement from the JUNO first data [JUNO1st]. In all the three panels, the improvement is quite sizable which clearly shows the importance of medium baseline reactor experiments. In all three panels, the improvement is quite sizable, which clearly demonstrates the importance of medium baseline reactor experiments. Quantitatively, from the narrowing of the $\Delta\chi^{2}$ profiles, one can estimate that JUNO has reduced the uncertainties on these effective mass limits by approximately $22.0\%$ for $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$, $22.5\%$ for $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$, and $23.1\%$ for $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$. The corresponding $3\,\sigma$ intervals after including the JUNO first data are $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(16.36,\,21.48)\,$meV, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}=(1.268,\,1.764)\,$meV, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}=(3.484,\,3.964)\,$meV. Although JUNO does not measure the $0\nu 2\beta$ decay directly in the first several years, it plays an important role in reducing the uncertainty from particle physics side.

### v2

@@ 1 Introduction
The baryon asymmetry of the Universe, namely that almost no anti-matter exists in Nature [1, 2], is one of the main problems in contemporary astroparticle physics and cosmology. An inflationary phase at the beginning of our Universe dilutes all pre-existing matter [3] which requires a mechanism to understand the observed baryon asymmetry. The Sakharov conditions [4] summarize the essential ingredients which must be fulfilled. One very well motivated mechanism is leptogenesis [5] which is typically related to Majorana neutrino masses with additional CP phases emerging by the seesaw mechanism [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
Verifying the Majorana nature of neutrinos is therefore very important for justifying the vanilla leptogenesis mechanism. Although the ultra-violet physics at high energy is difficult to observe with current technology, with two exceptions of cosmic string [20] and cosmological collider [21], the nature inherited by light active neutrinos can manifest itself in the neutrinoless double beta decay ($0\nu 2\beta$) process where a mother nuclei decays into a daughter nuclei plus two electrons, $^{Z}_{A}N\rightarrow{}^{Z+2}_{A}N+2e^{-}$ [22, 23, 24, 25]. With two electrons appearing in the final state, such process indicates that the lepton number is violated by two. In the minimal scenario, the $0\nu 2\beta$ decay can be induced by an effective Majorana mass $m_{ee}$ of the light active neutrinos in the electron flavor. Reciprocally, a lepton number violating process would finally contribute to a light Majorana mass [26].
@@
where $G^{0\nu}$ is the phase space factor [27, 28, 29, 30] associated with the decay kinematics and atomic physics, $M^{0\nu}$ is the nuclear matrix element (NME), and $\langle m_{ee}\rangle$ is the effective electron neutrino mass from particle physics. Of these three parts, the phase space factor can be calculated very precisely while the NME suffers a lot from theoretical uncertainties [24]. The NME calculations involve complex nuclear many-body physics and depend on the nuclear structure models. There are multiple theoretical approaches such as the quasiparticle random phase approximation (QRPA) [31, 32, 33, 34, 35], interacting shell mode (ISM) [36, 37, 38, 39], energy density functional methods (EDF) [40, 41, 42], interacting boson model (IBM) [43, 44], and ab initio techniques [45, 46, 47, 48, 49, 50, 51]. Their predictions typically differ by a factor of $2\sim 3$, representing the main source of systematic uncertainty from the nuclear physics side [24].
On the other hand, the particle physics side has improved quite significantly. The improvement comes from neutrino oscillation experiments, beta decay spectrum measurement around the end point, and cosmology. Especially, there is complementarity between the reactor neutrino oscillation experiments and $0\nu 2\beta$ decay [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62] since both involves only the electron flavor and share the same set of neutrino oscillation parameters. To be exact, only two mixing angles, the reactor angle $\theta_{r}\equiv\theta_{13}$ and the solar angle $\theta_{a}\equiv\theta_{12}$ can appear in the effective Majorana mass $m_{ee}$ for $0\nu 2\beta$ decay and the disappearence probability $P_{ee}$ of reactor neutrino oscillation. In addition, the solar mass squared difference $\Delta m^{2}_{s}\equiv\Delta m^{2}_{21}$ and the atmospheric one $\Delta m^{2}_{a}\equiv\Delta m^{2}_{31}$ can also enter. It is interesting to see that these four oscillation parameters can be precisely measured with the combination of Daya Bay and JUNO [63, 64]. Especially, the recent global fits [65] and JUNO first data [66] have significantly reduced the uncertainty of the solar angle $\theta_{s}$. In addition, the KATRIN experiment provides the best measurement of the beta decay spectrum to put a $\sum_{i}m_{i}\leq 0.45$ eV limit at 90% C.L. (confidence level) [67, 68]. The most stringent bound on the absolute neutrino mass comes from the recent DESI BAO data [69, 70], plus other existing observations such as CMB [71].
@@ 2 Effective Majorana Mass for $0\nu 2\beta$ Decay
where $m_{i}$ are the neutrino mass eigenvalues and $U_{ei}$ are the PMNS mixing matrix elements. We follow the PMNS matrix parametrization [72],
@@
A particularly illuminating perspective emerges when treating $\langle m_{ee}\rangle$ as a vector sum in the complex plane which can give a geometrical view [73, 74, 58]. We can decompose it into three contributions,
@@
The behavior of $\langle m_{ee}\rangle$ significantly depends on the ordering of the neutrino mass, with fundamentally different phenomenological implications for each scenario. The main distinction arises from the lightest mass, $m_{1}$ for NO and $m_{3}$ for IO. This difference has direct consequences for the allowed range of $\langle m_{ee}\rangle$, particularly in the limit where the lightest mass is smaller than $0.1$ eV. Note that distinguishing the mass ordering is the major physics goal of the JUNO experiment [64].
@@
The interplay between the neutrino oscillation parameters and the predictions for neutrinoless double-beta decay has been extensively studied in the literature, with detailed analysis of how experimental uncertainties propagate into the allowed ranges for the effective Majorana mass $\langle m_{ee}\rangle$ [75, 52, 53, 54, 55, 57, 62]. Especially, the uncertainty in the solar angle $\theta_{s}$ has significant effect on the upper and lower limits of the effective mass $\langle m_{ee}\rangle$ [55, 57]. In this section, we perform an updated analysis of [57] by incorporating the latest global fit results from neutrino oscillation experiments [65] and more crucially the first measurement of solar oscillation parameters $\sin^{2}\theta_{s}$ and $\Delta m^{2}_{s}$ from the JUNO experiment [66]. Our goal is to quantify how the improved determination of the solar mixing angle $\theta_{s}$ from the JUNO initial 59.1-day dataset impacts the theoretical predictions of $\langle m_{ee}\rangle$.
@@
Although both mass squared differences are involved in the electron flavor effective mass $\langle m_{ee}\rangle$, the limits are dominated by the atmohsperic one $\Delta m^{2}_{a}$ with much larger value. Further, the upper limit $\langle m_{ee}\rangle^{\rm IO}_{\rm max}$ even becomes independent of the solar angle $\theta_{s}$. Since both the atmospheric mass squared difference $\Delta m^{2}_{a}$ and the reactor angle $\theta_{r}$ have been measured quite precisely with just roughly 1% error [65], there is almost no uncertainty in the upper limit for IO. This has been clearly shown in Fig. 1 as the filled regions for IO have roughly the same upper boundary.
@@
For comparison, the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ depends on not just $\Delta m^{2}_{a}$ and the reactor angle $\theta_{r}$ but also the solar angle $\theta_{s}$. Although $\theta_{s}$ has also been measured very precisely with only 2% uncertainty [65], the combination $\cos 2\theta_{s}$ that appears in $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ can vary by 1/3 within the $3\,\sigma$ range. Even in the left panel of Fig. 1 with logarithmic scale, the effect of improving the solar angle measurement with JUNO can be clearly seen. Since the lower limit $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$ corresponds to the maximal value of half-life time $T^{0\nu}_{1/2}$ according to Eq. (1.1), such uncertainty would have significant consequences when interpreting the experimental data. Any improvement in the determination of $\theta_{s}$ would significantly reduce the uncertainty on the lower bound of $\langle m_{ee}\rangle$ for IO [75, 57].
@@
These two limits contain two mass squared differences and two mixing angles. Since the reactor angle $\theta_{r}$ is not large, its consine is roughly $c_{r}\approx 1$. The relative size between the two terms in Eq. (3.3) is then determined by comparing the ratios $\Delta m^{2}_{s}/\Delta m^{2}_{a}$ and $s^{2}_{r}/s^{2}_{s}$. According to the global fit [65], $\sqrt{\Delta m^{2}_{s}/\Delta m^{2}_{a}}\approx 0.17$ and $s^{2}_{r}/s^{2}_{s}\approx 0.07$ which means the first term $\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}$ dominates. Note that the solar angle $s^{2}_{s}$ appears in the first term. The current $3\,\sigma$ variation of $s^{2}_{s}$ can be as large as 23% [65]. Although the solar mass squared difference $\Delta m^{2}_{s}$ itself varies by $\sim 15\%$ at $3\,\sigma$, its contribution to $\langle m_{ee}\rangle$ is only $\sim 7.5\%$. So $\theta_{s}$ is the dominant source of uncertainty for NO. With all factors combined, the total effect of those uncertainties from the current data can be clearly seen in the orange bands of the left and right panels of Fig. 1. The difference is much more visible for the lower limits. This is because the variation in the first term $\sqrt{\Delta m^{2}_{s}}c^{2}_{r}s^{2}_{s}$ actually affects the central values as well and the whole NO branch in the limit of vanishing lightest mass moves vertically. While the lower limit $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$ is different between the two terms in Eq. (3.3), the upper one $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$ is a sum with larger value. Then the relative variation is much smaller for the upper limit.
@@
To make the influence of $\theta_{s}$ on the effective mass $\langle m_{ee}\rangle$ more visible, we have included two linear scale plots for IO and NO as the middle and right panels of Fig. 1. From this figure, it is also possible to infer from the orange bands both the minimum and maximum values of the effective mass for each mass ordering in the vanishing lightest mass scenario. Explicitly, their $3\,\sigma$ intervals are $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(15.72\sim 22.28)\,\text{meV}$, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}=(1.188\sim 1.828)\,\text{meV}$, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}=(3.404\sim 4.028)\,\text{meV}$, resulting from the propagation of the oscillation parameter uncertainties. While these uncertainties appear modest in $\langle m_{ee}\rangle$, their influence on the $0\nu 2\beta$ experiments is much more significant [57]. Especially, the half-life time scales inversely with the effective mass squared, $T^{0\nu}_{1/2}\propto 1/|\langle m_{ee}\rangle|^{2}$ according to Eq. (1.1). The variation in $\langle m_{ee}\rangle$ would get amplified when propagating to the half-life time. Further, to reach the required sensitivity on $T^{0\nu}_{1/2}$ the expected experimental exposure scales [76], $M\times t\propto(T^{0\nu}_{1/2})^{2}$ where $M$ is the detector target mass and $t$ is the running time, with the half-life time squared. Finally, the experimental exposure can differ by a factor of 4.3 for IO and 2.5 for NO. The improvement on the uncertainty from the particle physics side by JUNO is expected to significantly reduce the uncertainty in the experimental design and data interpretation.
@@
Following the first data release from JUNO, which reported 59.1 days of data taking [66], the precision on the solar angle $\theta_{s}$ and the solar mass splitting $\Delta m^{2}_{s}$ has been significantly improved to $\sin^{2}\theta_{s}=0.3092\pm 0.0087$ and $\Delta m^{2}_{s}=(7.50\pm 0.12)\times 10^{-5}\,$eV2. This corresponds to a reduction of uncertainty by approximately a factor of $0.741$ at $3\,\sigma$ than the global fit results from NuFIT 6.0 [65]. According to Eq. (3.2), this translates to around 22% reduction in the uncertainty of $\langle m_{ee}\rangle^{\rm IO}_{\rm min}\propto c^{2}_{s}-s^{2}_{s}$ at $3\,\sigma$. This improvement was achieved through JUNO’s unique capability to observe the low-frequency mode of the reactor antineutrino oscillation [63],
@@
For comparison, Fig. 1 shows four uncertainty scenarios. First, the theoretical prediction with perfect knowledge of oscillation parameters (red) that represents the ultimate goal of mitigating the uncertainty from the particle physics side to a negligible extend [57]. Second, the current $3\,\sigma$ range with the pre-JUNO global fits (orange) that shows what other experiments have achieved altogether [65]. Between these two cases, there is a huge space which is exactly the existing uncertainty from particle physics. In the middle, we have show two uncertainty scenario from the JUNO experiment. The blue band shows the reduced uncertainty with the JUNO first data release [66]. The comparison between the orange (pre-JUNO) and blue (post-JUNO) bands clearly illustrates the impact of the improved solar angle measurement by JUNO with a noticeable reduction in the width of the allowed regions for both mass orderings. The green one corresponds to the expected sensitivity with 6 years of data collection [64].
@@
For comparison, we have also shown those expected sensitivities from $0\nu 2\beta$ decay experiments as horizontal lines. The most stringent current bounds come from KamLAND-Zen (136Xe), yielding $T_{1/2}^{0\nu}>2.3\times 10^{26}$ yr at $90\%$ C.L., which translates to $\langle m_{ee}\rangle<(36\sim 156)$ meV depending on the nuclear matrix elements (NME) [77, 78]. This bound is indicated as the hashed region with solid boundaries in the left and middle panels of Fig. 1. The current experiments are beginning to probe the IO parameter space, particularly in the region where $m_{\text{lightest}}\gtrsim 10$ meV. The next-generation $0\nu\beta\beta$ experiments, including nEXO [79, 80], LEGEND-1000 [81, 82], and CUPID [83, 84], are projected to reach sensitivities of $\langle m_{ee}\rangle<(5\sim 12)$ meV, which would fully cover the IO prediction band and begin to probe the NO region.
@@
Fig. 2: One-dimensional $\Delta\chi^{2}$ for the $0\nu 2\beta$ effective mas limits $\langle m_{ee}\rangle^{\text{IO}}_{\text{min}}$ (left), $\langle m_{ee}\rangle^{\text{NO}}_{\text{min}}$ (center), and $\langle m_{ee}\rangle^{\text{NO}}_{\text{max}}$ (right) according to the neutrino oscillation global fit (blue) [65] or the combined results with both global fit and JUNO (red) [66].
@@
To make the influence of the JUNO first data [66] on the $0\nu 2\beta$ effective mass limits more explicit, Fig. 2 shows the one-dimensional $\Delta\chi^{2}$ profiles for $\langle m_{ee}\rangle^{\text{IO}}_{\text{min}}$, $\langle m_{ee}\rangle^{\text{NO}}_{\text{min}}$, and $\langle m_{ee}\rangle^{\text{NO}}_{\text{max}}$ with vanishing lightest mass ($m_{\text{lightest}}\to 0$). In this limit, the residual uncertainty is contributed by the oscillation parameters $\Delta m^{2}_{a}$, $\Delta m^{2}_{s}$, $\theta_{s}$, and $\theta_{r}$, as discussed before. The gray dashed horizontal lines indicate the $1\,\sigma$, $2\,\sigma$, and $3\,\sigma$ confidence levels ($\Delta\chi^{2}=1$, $4$, and $9$, respectively). For comparison, the blue curves represent the NuFIT 6.0 global fit of neutrino oscillation experiments [65] while the red curves show the improvement from the JUNO first data [66]. In all the three panels, the improvement is quite sizable which clearly shows the importance of medium baseline reactor experiments. In all three panels, the improvement is quite sizable, which clearly demonstrates the importance of medium baseline reactor experiments. Quantitatively, from the narrowing of the $\Delta\chi^{2}$ profiles, one can estimate that JUNO has reduced the uncertainties on these effective mass limits by approximately $22.0\%$ for $\langle m_{ee}\rangle^{\rm IO}_{\rm min}$, $22.5\%$ for $\langle m_{ee}\rangle^{\rm NO}_{\rm min}$, and $23.1\%$ for $\langle m_{ee}\rangle^{\rm NO}_{\rm max}$. The corresponding $3\,\sigma$ intervals after including the JUNO first data are $\langle m_{ee}\rangle^{\rm IO}_{\rm min}=(16.36,\,21.48)\,$meV, $\langle m_{ee}\rangle^{\rm NO}_{\rm min}=(1.268,\,1.764)\,$meV, and $\langle m_{ee}\rangle^{\rm NO}_{\rm max}=(3.484,\,3.964)\,$meV. Although JUNO does not measure the $0\nu 2\beta$ decay directly in the first several years, it plays an important role in reducing the uncertainty from particle physics side.